OpenStudy (anonymous):

what's 4-(16/3)....i hate fraction

OpenStudy (anonymous):

-4/3

OpenStudy (anonymous):

please show steps thanks in advance

OpenStudy (anonymous):

4 is nothing but 12/3 so 4 -16/3 = 12/3-16/3 = -4/3

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

you are welcome